# QCD phase-transition under the light of Thermofractal

### v1

@@ III The gluonic phase
Figure 1: Temperature dependence of the Polyakov loop $\langle L\rangle$ in $SU(3)$ gauge theory. The black circles represent $(2+1)$-flavor lattice QCD data [2]. The solid blue line denotes the Two-Phase Thermofractal model, which incorporates a gluonic phase ($q=12/11$) below $T_{c}$ and a quark phase ($q=8/7$) above $T_{c}$. The solid green line represents the rigid Gross-Witten-Wadia (GWW) transition ($q=1$), which exhibits a sharp third-order discontinuity at $T_{c}$. The dashed red line indicates the pure quark phase ($q=8/7$) extrapolated to the vacuum, highlighting the necessity of the gluonic phase at low temperatures. The best fit parameters are: $\Lambda=0.141\pm 0.002$ and $L_{c}=0.201\pm 0.005$. The reduced $\chi^{2}$ is 1.12. Cross-validation with HotQCD data [1] (red symbols, not used in calibration) yields similar agreement, confirming model independence.
@@ IV Analysis and Discussion
The results are reported in Fig. 1, where it is possible to observe that the theoretical curve for $N_{f}=3$ accurately describes the data in the entire range of temperatures. Note that the continuation of a pure $N_{f}=3$ regime (dashed-green line) cannot completely describe the data below $T_{c}$. Only after the introduction of the gluonic field can the full lattice data be accurately described in the full range of temperatures. This aspect shows that the gluonic regime is crucial for an accurate description of the lattice data. While our two-phase model captures the crossover, finite-volume effects in lattice data may slightly bias low-T tails [4], which probably can be accounted for by a recalibration of the parameter $L_{c}$.
@@
The physical significance of $\sigma^{2}(T)$ lies in the determination of the quantum fluctuations at different hierarchical levels of the thermofractal. Effectively, it works as a smoothing parameter for the eigenvalue density $\rho(\theta)$. In the limit $N_{f}\rightarrow\infty$, the transition becomes sharper, and the results approach the GWW line. In fact, this situation corresponds to $q\rightarrow 1$, when the thermofractal formalism recovers the GWW, and the lack of the temperature-dependent scaling results in a rigid potential that forces a hard gap. In our model, $\sigma^{2}(T)$ represents the thermal noise within the fractal clusters; as $T$ increases, this noise overcomes the Vandermonde repulsion, allowing eigenvalues to occupy the entire colour circle $\pi,\pi]$ without a non-analytic jump. This mechanism reconciles the rigid matrix model topology with the continuous crossover observed in $(2+1)$-flavour lattice QCD.

### v2

@@ III The gluonic phase
Figure 1: Temperature dependence of the Polyakov loop $\langle L\rangle$ in $SU(3)$ gauge theory. The blue squares represent $(2+1)$-flavor lattice QCD data [2]. The solid black line denotes the Two-Phase Thermofractal model, which incorporates a gluonic phase ($q=12/11$) below $T_{c}$ and a quark phase ($q=10/9$) above $T_{c}$, assuming $N_{f}=3$. The dotted black line represents the rigid Gross-Witten-Wadia (GWW) transition ($q=1$), which exhibits a sharp third-order discontinuity at $T_{c}$. The dashed black line indicates the pure quark phase ($q=10/9$) extrapolated to temperatures below $T_{c}$, highlighting the necessity of the gluonic phase at low temperatures. The dashed blue line indicates the QCD result ($N_{f}=6$), and the dashed red line shows an extrapolation of the pure gluonic phase to $T>T_{c}$. All cases collapse to the gluonic regime for $T<T_{c}$. The best fit parameters are: $\Lambda=0.141\pm 0.002$ and $L_{c}=0.201\pm 0.005$. The reduced $\chi^{2}$ is 1.12. Cross-validation with HotQCD data [1] (red symbols, not used in calibration) yields similar agreement, confirming model independence.
@@ IV Analysis and Discussion
The results are reported in Fig. 1, where it is possible to observe that the theoretical curve for $N_{f}=3$ accurately describes the data in the entire range of temperatures. Note that the continuation of a pure $N_{f}=3$ regime cannot completely describe the data below $T_{c}$. Only after the introduction of the gluonic field can the full lattice data be accurately described in the full range of temperatures. This aspect shows that the gluonic regime is crucial for an accurate description of the lattice data. While our two-phase model captures the crossover, finite-volume effects in lattice data may slightly bias low-T tails [4], which probably can be accounted for by a recalibration of the parameter $L_{c}$.
@@
The physical significance of $\sigma^{2}(T)$ lies in the determination of the quantum fluctuations at different hierarchical levels of the thermofractal. Effectively, it works as a smoothing parameter for the eigenvalue density $\rho(\theta)$. In the limit $N_{f}\rightarrow\infty$, the transition becomes sharper, and the results approach the GWW result. In fact, this situation corresponds to $q\rightarrow 1$, when the thermofractal formalism recovers the GWW, and the lack of the temperature-dependent scaling results in a rigid potential that forces a hard gap. In our model, $\sigma^{2}(T)$ represents the thermal noise within the fractal clusters; as $T$ increases, this noise overcomes the Vandermonde repulsion, allowing eigenvalues to occupy the entire colour circle $\pi,\pi]$ without a non-analytic jump. This mechanism reconciles the rigid matrix model topology with the continuous crossover observed in $(2+1)$-flavour lattice QCD.
@@ V Conclusion
VI Acknowledgements
This work was funded by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq-Brazil), grant number 306093/2022-7, and FAPESP, grant number 2024/01533-7.